Circular No. 7009

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1550–564 J. Orosz, Pennsylvania State University; and C. Bailyn and R. Jain,

Yale University, report YALO consortium observations using the Yale 1-m telescope (+ ANDICAM CCD camera) at the Cerro Tololo Interamerican Observatory: “We have identified a possible optical counterpart to the recent x-ray transient XTE J1550–564 (IAUC 7008) in V -band images obtained on Sept. 8.99 UT. The candidate is located at α = 15^{h}50^{m}58^{s}.78, δ = −56^{o}28′35′′.0 (equinox 2000.0), with estimated errors on each value of ∼ 1′′–2′′. This object is not visible on an image obtained from the Digitized Sky Survey. We estimate a magnitude of V = 16, based on comparions with nearby stars selected from the Hubble Guide Star Catalog. Further observations, especially spectroscopic confirmation, are urgently needed. A finding chart may be obtained via the World Wide Web at

http://www.astro.psu.edu/users/orosz/xte.html.”

Corrigendum. On IAUC 7008, line 15, for 6 × 10^{−8} read 6 × 10^{−9}

AX J0103–722 = 1SAX J0103.2–7209 J. Yokogawa and K. Koyama, Kyoto University, report: “An ASCA

observation made on 1996 May 21–23 revealed quasisinusoidal pulsations at a period of 348.9 ± 0.3 s from a source at α = 1^{h}03^{m}15^{s}, δ = −72^{o}09′01′′ (equinox 2000.0; ± 1′ at 90-percent confidence), designated as AX J0103– 722. The spectrum was well fitted with a power law with photon index 0.8 ± 0.2 and an absorbed flux of 1.0 × 10^{−12} erg s^{−1} cm^{−2} (2–10 keV). AX J0103–722 coincides in position with the new pulsar 1SAX J0103.2–7209 (period 345.2 ± 0.1 s; IAUC 6999). Thus the period derivative (dP /dt) is found to be −1.7 s/yr. This source was also detected on 1993 May 12 and 1997 Nov. 14 with respective fluxes of 1.2 × and 1.1 × 10^{−12} erg s^{−1} cm^{−2} (2–10 keV); however, no pulsation was discovered.”

R CORONAE BOREALIS This variable star appears to be fading, as indicated by the following

visual magnitude estimates: Aug. 18.87 UT, 6.2 (J. Shanklin, Cambridge, England); 23.92, 6.1 (A. Baransky, Kiev, Ukraine); 24.81, 6.6 (S. Foglia, Milan, Italy); 24.94, 6.7 (Shanklin); 27.81, 7.0 (Foglia); Sept. 4.81, 8.1 (Baransky); 5.86, 8.1 (A. Pereira, Cabo da Roca, Portugal); 6.84, 8.4 (Baransky); 7.88, 8.3 (Pereira); 8.86, 8.2 (C. Vitorino, Cabo da Roca, Portugal).

1998 September 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green